# Full Counting Statistics of Yu-Shiba-Rusinov Bound States

### v2

@@ II Full Counting Statistics: A Reminder
[…]
@@
[…]
@@ III Keldysh Action
where $\hat{G}(\chi)$ contains the information of the reservoir Green’s functions (GFs) and $\hat{S}$ is the normal-state scattering matrix whose structure will be explained in what follows. First, $\hat{G}(\chi)$ is a GF in the lead-time-Keldysh-spin-Nambu space, denoted by the symbol $(\ \hat{}\ )$. In this work, we focus on two-terminal settings for which this GF adopts the generic form
@@
which is a block-diagonal matrix with the two time-Keldysh-spin-Nambu GFs as entries on the diagonal. The counting field, similarly to the voltage, can be gauged away from the right onto the left lead. In what follows the right terminal will always be superconducting $G_{\mathrm{R}}=G_{\mathrm{SC}}$, while the left terminal can be either normal or superconducting. The GFs are in fact infinite matrices in time space, which means that the element $(t,t^{\prime})$ of the GF $G_{\mathrm{L,R}}(\chi)$ is the Keldysh GF $\check{g}_{\rm L,R}(\chi,t,t^{\prime})$. In other words, the lead GF is an infinitely large matrix in time space whose entries are $16\times 16$-matrices in lead-Keldysh-spin-Nambu space. For a detailed explanation of the structure of the GFs we refer to Appendix A. Equivalently, the normal state scattering matrix $\hat{S}$ is also an infinite matrix in time space whose $(t,t^{\prime})$ entry is a $16\times 16$-matrix in lead-Keldysh-spin-Nambu space, which we denote as $\tilde{s}(t,t^{\prime})$, where the symbol $(\ \tilde{}\ )$ indicates that the scattering matrix is expressed in lead-Keldysh-spin-Nambu space. However, the scattering matrix only depends on the relative time $t_{\mathrm{rel}}=t-t^{\prime}$, meaning that $\tilde{s}(t,t^{\prime})=\tilde{s}(t_{\mathrm{rel}})$. To elaborate on the structure of the scattering matrix, in a two-terminal setting it can be written in lead space as follows
@@
[…]
@@
[…]
@@ IV Single-impurity YSR junctions: Model and scattering matrix
$\bar{g}^{\rm r/a}_{j}(E)$ $=$ $\sum_{\bm{k}}\bar{g}_{\bm{k},j}^{\rm r/a}(E)$ (26)
$=$ $\sigma_{0}\otimes\frac{-1}{\sqrt{\Delta_{j}^{2}-(E\pm\imath\eta_{j})^{2}}}\begin{pmatrix}E\pm\imath\eta_{j}&\Delta_{j}e^{\imath\phi_{j}}\\ \Delta_{j}e^{-\imath\phi_{j}}&E\pm\imath\eta_{j}\end{pmatrix},$ (29)
@@
where the index ()$({}_{\rm e})$ refers to extracting the first and third component of the respective quantity in spin-Nambu space. The dressed electron impurity retarded GF can be calculated using the self-energies as follows
@@ V Single-impurity YSR junctions: Normal conducting tip
[…]
where $P_{n}^{\sigma}(E,V)$ corresponds to the probability of transferring $n$ charges across the junction for spin $\sigma$. The single-quasiparticle tunneling probabilities from tip to substrate ($n=1$) and from substrate to tip ($n=-1$) are given by
@@
while the corresponding Andreev reflection probabilities from tip to substrate ($n=2$) and from substrate to tip ($n=-2$) are given by
@@
where $-f^{\rm a}f^{\rm r}$ corresponds to the energy-dependent Cooper pair density. The other contributions are zero ($P_{-1/-2}^{\sigma}=0$) as there is no current flowing to the normal metal without thermal excitation. For negative voltages, we obtain
$P_{-1}^{\sigma}(E,V<0)$ $=$ $\frac{1}{D^{\sigma}}\left\{\begin{array}[]{cc}\rho_{\mathrm{S}}|t_{\rm h}^{\sigma}|^{2}+|t_{\rm h}^{\sigma}|^{2}|t_{\rm e}^{\sigma}|^{2}(1+f^{\rm a}f^{\rm r}-g^{\rm a}g^{\rm r}-2\rho_{\mathrm{S}})&\mbox{if}\;E\in[-e|V|,0]\\ \rho_{\mathrm{S}}|t_{\rm e}^{\sigma}|^{2}+|t_{\rm e}^{\sigma}|^{2}|t_{\rm h}^{\sigma}|^{2}(1+f^{\rm a}f^{\rm r}-g^{\rm a}g^{\rm r}-2\rho_{\mathrm{S}})&\mbox{if}\;E\in[0,e|V|]\\ 0&\,\textrm{otherwise}\\ \end{array}\right.$ (63)
@@
$S(V)$ $=$ $\frac{e^{2}}{h}\sum_{\sigma=\Uparrow,\Downarrow}\int^{\infty}_{-\infty}dE\left\{\sum^{2}_{n=-2}n^{2}P_{n}^{\sigma}(E,V)-\right.$
@@
[…]
@@ VI Single-impurity YSR junctions: Comparison with shot noise measurements
Very recently, Thupakula and coworkers reported the first measurements of the shot noise through a magnetic impurity on a superconducting substrate featuring YSR states [49]. To be precise, these researchers employed shot-noise scanning tunneling microscopy to measure the nonequilibrium current fluctuations through a magnetic impurity deposited on ${}_{2}$ at temperatures around $0.7$ K, way below the critical temperature of this superconductor. The STM tip was made of a normal metal and the main observation was the appearance of a Fano factor above 1. This fact was interpreted as the evidence of the contribution of an Andreev reflection to the charge transport. Moreover, those authors presented a theoretical analysis that suggested that the broadening of the YSR state was of the order of $1$ $\mu$eV, which was clearly below the thermal energy ($k_{\rm B}T$) in this experiment demonstrating that shot noise can probe energy scales that are not accessible in conventional tunneling spectroscopy measurements. The goal of this section is to provide a thorough analysis of those experimental results in the light of the theory presented in Sec. V.
@@
Figure 7: Conductance (a) and Fano factor (b) as a function of voltage. The symbols correspond to the experimental data of Ref. [49] that were obtained with a normal tip at $T=0.7$ K and a magnetic impurity deposited on a ${}_{2}$ superconducting surface. The orange lines in both panels correspond to a fit with our theory using a two-channel model, as described in the main text. The yellow lines in both panels correspond to the theory results considering only the contribution of the magnetic channel (see explanation in the text).
These experimental results were analyzed in Ref. [49] in the light of a model for the YSR state similar in spirit to ours, but just focusing on voltages inside the gap. To be precise, these authors used an approximation similar to that summarized in our Eqs. (69) and (70). Although such a simplified model qualitatively captures the physics of the YSR states, it is obvious that it cannot provide an overall correct picture of the experimental results. Actually, the major problem is that no model based on a single magnetic channel can describe the simultaneous appearance in the conductance of YSR peaks and pronounced coherent peaks as those in Fig. 7(a). As we explained above, these models do not predict the appearance of the coherent peaks simply due to the conservation of the number of states. So, as a way out to provide an overall consistent fit of the experimental results, we propose that there are at least two channels or pathways for the current: one due to the magnetic impurity and another non-magnetic channel, which probably results from the direct tunneling of the STM tip to the superconducting substrate. This is actually a solution that has been proposed before to explain the reported conductance spectra, see for instance Refs. [23, 27]. There is another subtlety that we need to take into account, namely ${}_{2}$ is not a regular BCS superconductor. In fact, it has been shown by several groups that 2H-NbSe2 can be described as a two-band superconductor [67, 68, 20]. So, we propose here to explain the experimental results assuming that the transport takes place via two independent channels, one magnetic that is described by our YSR model of the previous sections and a non-magnetic channel that proceeds from the normal tip to the ${}_{2}$ substrate that we describe as a two-band superconductor. Moreover, we shall assume that this non-magnetic channel can be described in the tunneling regime, i.e., taking only into account the quasiparticle tunneling at the lowest order in transmission. We proceed now to describe the details of such a two-channel model.
First, for the non-magnetic channel we describe the superconductivity in ${}_{2}$ using the two-band model described in Ref. [68]. In this model one considers that the SC order parameter has two energy-dependent components $\Delta_{1}(E)$ and $\Delta_{2}(E)$ that can be computed by solving the following self-consistent set of equations
@@
[…]
@@
Something that is very important to emphasize is the fact that, as we showed in Sec. V, the Fano factor is very sensitive to the broadening or lifetime of the YSR states. In our fit we obtained a value of $\eta=1.1$ $\mu$eV for the Dynes’ parameter in the superconducting substrate. Using Eq. (72) and the rest of the parameter values extracted from the fit, we obtain that $W\approx\eta=1.1$ $\mu$eV, which is much smaller than the thermal energy in this case ($k_{\rm B}T\approx 60$ $\mu$eV). Thus, as stated in Ref. [49], the analysis of the noise and the corresponding Fano factor allows us to have access to energy and time scales that are usually out of the scope of conventional conductance measurements due to thermal broadening.
@@ VII Single-impurity YSR junctions: Superconducting tip
where $P_{n}^{\sigma}(E,V)$ corresponds to the probability of transferring $n$ charges across the junction for spin $\sigma$ and $E$ is the Floquet energy. Notice that the main difference with respect to Eq. (47) is the fact that $n$ now runs from $\infty$- ∞ to $\infty$ due to the occurrence of MARs and we integrate over a finite energy range. Again, from the knowledge of the probabilities $P_{n}^{\sigma}(E,V)$ we can easily compute the current and the noise, which are given by the standard formulas of a multinomial distribution
[…]
$S(V)$ $=$ $\frac{e^{2}}{h}\sum_{\sigma=\Uparrow,\Downarrow}\int^{eV}_{-eV}dE\left\{\sum^{\infty}_{n=-\infty}n^{2}P_{n}^{\sigma}(E,V)-\right.$
@@ VIII.1 Model and scattering matrix
¯$\bar{\mathcal{H}}_{\bm{k},j}$ $=\sigma_{0}\otimes\left(\xi_{\bm{k},j}\tau_{3}+\Delta_{j}\tau_{1}\right)$ (100)
¯$\bar{\mathcal{H}}_{\mathrm{imp},j}$ $=U_{j}\sigma_{0}\otimes\tau_{3}+J_{j}(\sigma_{3}\otimes\tau_{0})$ (101)
@@ Appendix A Keldysh Green function formalism
$\bar{g}^{\rm r/a}(E)$ $=\sigma_{0}\otimes\frac{-\imath}{\sqrt{\Delta^{2}-(E\pm\imath\eta)^{2}}}\begin{pmatrix}(E\pm\imath\eta)&\Delta e^{\imath\phi}\\ -\Delta e^{-\imath\phi}&-(E\pm\imath\eta)\end{pmatrix}$ (147)
@@
where we have used the spinor basis $(\Psi_{\uparrow}^{\dagger},\Psi_{\downarrow},\Psi_{\downarrow}^{\dagger},-\Psi_{\dagger})$ and the Pauli matrices $\sigma_{i}$ in spin space, $\sigma_{0}$ being the identity. Note that we already eliminated the $k$-dependence by summing it out. The bar ($\ \bar{}$) indicates that a quantity is expressed in spin-Nambu space. The SC is described by a gap $\Delta>0$ and the phase of the SC condensate $\phi$. The regularization parameter $\eta$ should in principle tend to zero in these expressions, but it can be kept constant to describe in simple terms inelastic mechanisms that tend to broaden the electronic states (Dynes’ parameter). Note that we are using here GFs subjected to the normalization condition $\bar{g}^{\rm r/a}\bar{g}^{\rm r/a}=\bar{1}$, where $\bar{1}$ is the identity in spin-Nambu space. Notice also that the GFs are block-diagonal in spin space
@@
with the retarded and advanced GF in spin-Nambu basis from Eq. (147) and the Keldysh component of the GF $\bar{g}^{\rm k}=(\bar{g}^{\rm r}-\bar{g}^{\rm a})(1-2f)$ with the Fermi function of the superconductor $f(E)=1/(e^{E/k_{\mathrm{B}}T}+1)$ with the temperature $T$ and the Boltzman constant $k_{\mathrm{B}}$. The index $(\ \check{}\ )$ indicates that the GF is expressed in Keldysh-spin-Nambu space. The Keldysh GF can be brought into the time-domain by a Fourier transformation as follows:
@@
where $t_{\mathrm{rel}}=t-t^{\prime}$ is the relative time. If a reservoir is voltage biased, the voltage can be gauged onto the GF. Namely, for a constant bias voltage $V$, the SC phase is given by $\phi(t)=2eVt/\hbar$ (where the dc phase $\phi_{0}$ is already included in the GF in Eq. (147)). The voltage dependent GF is then obtained via
@@
[…]
@@
The other ingredient for the action in Eq. (9) is the normal state scattering matrix. The structure of the scattering matrix is explained in detail in Sec. IV for the single-impurity and in Sec. VIII for the double-impurity case. The scattering matrix is a function of the energy $E$. Thus, upon a Fourier transformation, the scattering matrix only depends on relative time $\tilde{s}=\tilde{s}(t_{\rm rel})$ where ($\ \tilde{\ }$) signifies that the quantity is expressed in lead-Keldysh-spin-Nambu space. Thus, its Wigner representation is trivial and its Floquet matrix $\breve{\mathcal{S}}$ is diagonal in Floquet space. In particular, the element $(n,m)$ of the Floquet matrix is given by

### v3

@@ II Full Counting Statistics: A Reminder
This is a binomial distribution where we define the number of attempts $M=\lceil eVt_{0}/h\rceil$, where $\lceil\ \rceil$ describes the next highest integer. This is justified by the fact that $t_{0}$ is chosen to be sufficiently large. Hence, it is evident that the transport can be described as $N$ particles being transmitted in individual tunneling processes transferring a single electron charge with a probability $\tau$.
@@
$\mathcal{A}_{t_{0}}(\chi)=\frac{t_{0}}{2h}\sum_{\sigma}\int dE\ln\left[\sum^{\infty}_{n=-\infty}P_{n}^{\sigma}(E,V)e^{\imath n\chi}\right],$ (8)
@@ III Keldysh Action
where $\hat{G}(\chi)$ contains the information of the reservoir Green’s functions (GFs) and $\hat{S}$ is the normal-state scattering matrix whose structure will be explained in what follows. First, $\hat{G}(\chi)$ is a GF in the lead-time-Keldysh-spin-Nambu space, denoted by the symbol (^)(\ $\hat{}\$). In this work, we focus on two-terminal settings for which this GF adopts the generic form
@@
which is a block-diagonal matrix with the two time-Keldysh-spin-Nambu GFs as entries on the diagonal. The counting field, similarly to the voltage, can be gauged away from the right onto the left lead. In what follows the right terminal will always be superconducting $G_{\mathrm{R}}=G_{\mathrm{SC}}$, while the left terminal can be either normal or superconducting. The GFs are in fact infinite matrices in time space, which means that the element $(t,t^{\prime})$ of the GF $G_{\mathrm{L,R}}(\chi)$ is the Keldysh GF $\check{g}_{\rm L,R}(\chi,t,t^{\prime})$. In other words, the lead GF is an infinitely large matrix in time space whose entries are $16\times 16$-matrices in lead-Keldysh-spin-Nambu space. For a detailed explanation of the structure of the GFs we refer to Appendix A. Equivalently, the normal state scattering matrix $\hat{S}$ is also an infinite matrix in time space whose $(t,t^{\prime})$ entry is a $16\times 16$-matrix in lead-Keldysh-spin-Nambu space, which we denote as $\tilde{s}(t,t^{\prime})$, where the symbol (~)(\ $\tilde{}\$) indicates that the scattering matrix is expressed in lead-Keldysh-spin-Nambu space. However, the scattering matrix only depends on the relative time $t_{\mathrm{rel}}=t-t^{\prime}$, meaning that $\tilde{s}(t,t^{\prime})=\tilde{s}(t_{\mathrm{rel}})$. To elaborate on the structure of the scattering matrix, in a two-terminal setting it can be written in lead space as follows
@@
$\mathcal{A}(\chi)=\frac{1}{2}\int_{-eV}^{eV}dE\,\mathrm{Tr}\left(\ln\left[\breve{\cal Q}(\chi,E)\right]-\ln\left[\breve{\cal Q}(0,E)\right]\right).$ (14)
@@
$\mathcal{A}(\chi)$ $=$ $\frac{1}{2}\int_{-eV}^{eV}dE\,\left(\ln\det\left[\breve{\cal Q}(\chi,E)\right]-\ln\det\left[\breve{\cal Q}(0,E)\right]\right)$ (15)
$=$ $\frac{1}{2}\int_{-eV}^{eV}dE\,\ln\frac{\det\breve{\cal Q}(\chi,E)}{\det\breve{\cal Q}(0,E)}$
$=$ $\frac{1}{2}\sum_{\sigma=\Uparrow,\Downarrow}\int_{-eV}^{eV}dE\,\ln\frac{\det\breve{\cal Q}^{\sigma}(\chi,E)}{\det\breve{\cal Q}^{\sigma}(0,E)}$
$=$ $\frac{1}{2}\sum_{\sigma=\Uparrow,\Downarrow}\int_{-eV}^{eV}dE\,\ln\mathcal{P}^{\sigma}(\chi,E),$
@@ IV Single-impurity YSR junctions: Model and scattering matrix
$\bar{g}^{\rm r/a}_{j}(E)$ $=$ $\sum_{\bm{k}}\bar{g}_{\bm{k},j}^{\rm r/a}(E)$
$=$ $\sigma_{0}\otimes\frac{-1}{\sqrt{\Delta_{j}^{2}-(E\pm\imath\eta_{j})^{2}}}\begin{pmatrix}E\pm\imath\eta_{j}&\Delta_{j}e^{\imath\phi_{j}}\\ \Delta_{j}e^{-\imath\phi_{j}}&E\pm\imath\eta_{j}\end{pmatrix}$,
@@
[…]
@@ V Single-impurity YSR junctions: Normal conducting tip
$\mathcal{A}_{t_{0}}(\chi)=\frac{t_{0}}{2h}\sum_{\sigma=\Uparrow,\Downarrow}\int^{\infty}_{-\infty}dE\ln\left[\sum_{n=-2}^{2}P_{n}^{\sigma}(E,V)e^{\imath n\chi}\right],$ (44)
where $P_{n}^{\sigma}(E,V)$ corresponds to the probability of transferring $n$ charges across the junction for spin $\sigma$. The single-quasiparticle tunneling probabilities from tip to substrate ($n=1$) and from substrate to tip (n=1n=-1) are given by
@@
while the corresponding Andreev reflection probabilities from tip to substrate ($n=2$) and from substrate to tip (n=2n=-2) are given by
@@
where $f^{\rm a}f^{\rm r}$ corresponds to the energy-dependent Cooper pair density. The other contributions are zero ($P_{-1/-2}^{\sigma}=0$) as there is no current flowing to the normal metal without thermal excitation. For negative voltages, we obtain
[…]
@@
$I(V)$ $=$ $\frac{e}{2h}\sum_{\sigma=\Uparrow,\Downarrow}\int^{\infty}_{-\infty}dE\sum^{2}_{n=-2}nP_{n}^{\sigma}(E,V),$ (64)
@@
$I(V)$ $=$ $\frac{2e}{h}\int_{-\infty}^{\infty}dE\,P^{\rm YSR}_{2}(E)=\frac{2e}{h}\pi W,$ (70)
$S(V)$ $=$ $\frac{8e^{2}}{h}\int_{-\infty}^{\infty}dE\,P^{\rm YSR}_{2}(E)(1-P^{\rm YSR}_{2}(E))$ (71)
@@ VI Single-impurity YSR junctions: Comparison with shot noise measurements
[…]
@@
Figure 7: Conductance (a) and Fano factor (b) as a function of voltage. The symbols correspond to the experimental data of Ref. [49] that were obtained with a normal tip at $T=0.7$ K and a magnetic impurity deposited on a 2H-NbSe2 superconducting surface. The orange lines in both panels correspond to a fit with our theory using a two-channel model, as described in the main text. The yellow lines in both panels correspond to the theory results considering only the contribution of the magnetic channel (see explanation in the text).
These experimental results were analyzed in Ref. [49] in the light of a model for the YSR state similar in spirit to ours, but just focusing on voltages inside the gap. To be precise, these authors used an approximation similar to that summarized in our Eqs. (66) and (67). Although such a simplified model qualitatively captures the physics of the YSR states, it is obvious that it cannot provide an overall correct picture of the experimental results. Actually, the major problem is that no model based on a single magnetic channel can describe the simultaneous appearance in the conductance of YSR peaks and pronounced coherent peaks as those in Fig. 7(a). As we explained above, these models do not predict the appearance of the coherent peaks simply due to the conservation of the number of states. So, as a way out to provide an overall consistent fit of the experimental results, we propose that there are at least two channels or pathways for the current: one due to the magnetic impurity and another non-magnetic channel, which probably results from the direct tunneling of the STM tip to the superconducting substrate. This is actually a solution that has been proposed before to explain the reported conductance spectra, see for instance Refs. [23, 27]. There is another subtlety that we need to take into account, namely 2H-NbSe2 is not a regular BCS superconductor. In fact, it has been shown by several groups that 2H-NbSe2 can be described as a two-band superconductor [67, 68, 20]. So, we propose here to explain the experimental results assuming that the transport takes place via two independent channels, one magnetic that is described by our YSR model of the previous sections and a non-magnetic channel that proceeds from the normal tip to the 2H-NbSe2 substrate that we describe as a two-band superconductor. Moreover, we shall assume that this non-magnetic channel can be described in the tunneling regime, i.e., taking only into account the quasiparticle tunneling at the lowest order in transmission. We proceed now to describe the details of such a two-channel model.
First, for the non-magnetic channel we describe the superconductivity in 2H-NbSe2 using the two-band model described in Ref. [68]. In this model one considers that the SC order parameter has two energy-dependent components $\Delta_{1}(E)$ and $\Delta_{2}(E)$ that can be computed by solving the following self-consistent set of equations
@@
$\rho_{i}(E)=\rho_{i}(E_{\rm F})\int d\theta\,\Re\left(\frac{|E|}{\sqrt{(1+\alpha\cos(\theta))\Delta^{2}_{i}(E)-E^{2}}}\right),$ (74)
@@
[…]
@@ VII Single-impurity YSR junctions: Superconducting tip
$\mathcal{A}_{t_{0}}(\chi)=\frac{t_{0}}{2h}\sum_{\sigma=\Uparrow,\Downarrow}\int^{eV}_{-eV}dE\ln\left[\sum_{n=-\infty}^{\infty}P_{n}^{\sigma}(E,V)e^{\imath n\chi}\right],$ (78)
where $P_{n}^{\sigma}(E,V)$ corresponds to the probability of transferring $n$ charges across the junction for spin $\sigma$ and $E$ is the Floquet energy. Notice that the main difference with respect to Eq. (44) is the fact that $n$ now runs from -$\infty$ to $\infty$ due to the occurrence of MARs and we integrate over a finite energy range. Again, from the knowledge of the probabilities $P_{n}^{\sigma}(E,V)$ we can easily compute the current and the noise, which are given by the standard formulas of a multinomial distribution
$I(V)$ $=$ $\frac{e}{2h}\sum_{\sigma=\Uparrow,\Downarrow}\int^{eV}_{-eV}dE\sum^{\infty}_{n=-\infty}nP_{n}^{\sigma}(E,V),$ (79)
@@ VIII.1 Model and scattering matrix
$\bar{\mathcal{H}}_{\bm{k},j}$ $=\sigma_{0}\otimes\left(\xi_{\bm{k},j}\tau_{3}+\Delta_{j}\tau_{1}\right)$ (97)
$\bar{\mathcal{H}}_{\mathrm{imp},j}$ $=U_{j}\sigma_{0}\otimes\tau_{3}+J_{j}(\sigma_{3}\otimes\tau_{0})$ (98)
@@ VIII.2 Results
$\mathcal{A}_{t_{0}}(\chi)=\frac{t_{0}}{2h}\int^{eV}_{-eV}dE\ln\left[\sum_{n=-\infty}^{\infty}P_{n}(E,V)e^{\imath n\chi}\right],$ (114)
$I(V)$ $=$ $\frac{e}{2h}\int^{eV}_{-eV}dE\sum^{\infty}_{n=-\infty}nP_{n}(E,V),$ (115)
@@ Appendix A Keldysh Green function formalism
$\bar{g}^{\rm r/a}(E)$ $=\sigma_{0}\otimes\frac{-\imath}{\sqrt{\Delta^{2}-(E\pm\imath\eta)^{2}}}\begin{pmatrix}(E\pm\imath\eta)&\Delta e^{\imath\phi}\\ -\Delta e^{-\imath\phi}&-(E\pm\imath\eta)\end{pmatrix}$
@@
where we have used the spinor basis $(\Psi_{\uparrow}^{\dagger},\Psi_{\downarrow},\Psi_{\downarrow}^{\dagger},-\Psi_{\dagger})$ and the Pauli matrices $\sigma_{i}$ in spin space, $\sigma_{0}$ being the identity. Note that we already eliminated the $k$-dependence by summing it out. The bar (¯\ $\bar{}\$) indicates that a quantity is expressed in spin-Nambu space. The SC is described by a gap $\Delta>0$ and the phase of the SC condensate $\phi$. The regularization parameter $\eta$ should in principle tend to zero in these expressions, but it can be kept constant to describe in simple terms inelastic mechanisms that tend to broaden the electronic states (Dynes’ parameter). Note that we are using here GFs subjected to the normalization condition $\bar{g}^{\rm r/a}\bar{g}^{\rm r/a}=\bar{1}$, where $\bar{1}$ is the identity in spin-Nambu space. Notice also that the GFs are block-diagonal in spin space
@@
$\check{g}(t_{\mathrm{rel}})=\int dE\,\check{g}(E)e^{\imath Et_{\mathrm{rel}}},$ (132)
where $t_{\mathrm{rel}}=t-t^{\prime}$ is the relative time. If a reservoir is voltage biased, the voltage can be gauged onto the GF. Namely, for a constant bias voltage $V$, the SC phase is given by $\phi(t)=2eVt/\hbar$ (where the dc phase $\phi_{0}$ is already included in the GF in Eq. (A)). The voltage dependent GF is then obtained via
@@
$\check{g}^{n}(E)\equiv\int_{\mathbb{R}}dt_{\mathrm{rel}}\frac{1}{\tau}\int_{-\tau/2}^{\tau/2}dt_{\mathrm{av}}e^{\imath Et_{\mathrm{rel}}+\imath n\Omega t_{\mathrm{av}}}\check{g}(t_{\mathrm{rel}},t_{\mathrm{av}}),$ (135)
@@
The other ingredient for the action in Eq. (9) is the normal state scattering matrix. The structure of the scattering matrix is explained in detail in Sec. IV for the single-impurity and in Sec. VIII for the double-impurity case. The scattering matrix is a function of the energy $E$. Thus, upon a Fourier transformation, the scattering matrix only depends on relative time $\tilde{s}=\tilde{s}(t_{\rm rel})$ where (~\ $\tilde{\ }\$) signifies that the quantity is expressed in lead-Keldysh-spin-Nambu space. Thus, its Wigner representation is trivial and its Floquet matrix $\breve{\mathcal{S}}$ is diagonal in Floquet space. In particular, the element $(n,m)$ of the Floquet matrix is given by